A jumbo (i.e., negotiable) CD has a quoted yield of 4.66%. This is the yield conventionally quoted for this security. It has 76 days to maturity. What is the effective annual yield of this security? Note it makes sense to first calculate the bond equivalent yield. Answer in percent to three decimal places. Omit the percent sign.
Answer #1

Bond Equivalent Yield = Quoted yield * [365 / 360] = 4.66% * 1.0139 = 4.725%

EAY = [1 + (BEY/m)]m - 1; m = no. of compounding periods in a year

= [1 + {0.04725 / (365/76)}](365/76) - 1

= [1 + 0.0098]4.8026 - 1 = 1.04814 - 1 = 0.04814, or 4.814%